OpenStudy (anonymous):

Let f(x)=√6x and g(x) = x -3. What's the smallest number that is in the domain of fog?

OpenStudy (anonymous):

Let f(x) = x2-1, g(x) = 3x-1 Domain f = R Domain g = R fog (x) = f [g(x)] (range of g C domain of f ) = f(3x-1) = (3x-1)2 - 1 = 9x2 + 1 - 6x - 1 = 9x2 - 6x Domain fog = R

OpenStudy (anonymous):

hope i helped?

OpenStudy (anonymous):

Thanks, but you kind of lost me here.

OpenStudy (anonymous):

@Lilium

OpenStudy (anonymous):

yes?